# Incorporating Flexible Image Conditioning into Text-to-Video Diffusion Models without Training

### v1

@@ 4 FlexTI2V
The setting of the proposed generic text-image-to-video (TI2V) generation problem is illustrated in Fig. 2. The input includes a textual prompt $y$, $N$ images $\bm{x}=\{x^{(n)}\}_{n=0}^{N-1}$ as visual conditions, and the expected position index $\bm{p}=\{p^{(n)}\}_{n=0}^{N-1}$ of each image in the output video. The goal is to synthesize a video consisting of $M$ frames $\bm{v}=\{v^{(m)}\}_{m=0}^{M-1}$ with each condition image at the expected position (i.e., $v^{(p^{(n)})}=x^{(n)}$ where $p^{(n)}\in[0,M-1]$). The output video is supposed to show a reasonable transition across condition images following the textual prompt. Our problem setting unifies many existing TI2V tasks in one formulation, such as image animation ($N=1$, $\bm{p}=\{0\}$), video rewinding ($N=1$, $\bm{p}=\{M-1\}$) and video frame interpolation ($N=2$, $\bm{p}=\{0,M-1\}$).
The major challenge of training-free TI2V generation is harmonizing the distinct denoising directions controlled by texts and images solely relying on frozen model parameters. Recent research [10] validates that global image information can be carried by and inferred from a portion of local patches. Inspired by this finding, the key insight of our FlexTI2V method is contextualizing each synthetic video frame with visual conditioning, by randomly swapping a portion of patches between video frames and condition images in the latent space at early denoising steps. To balance the fidelity to condition images and creativity in motion generation, we use a dynamic control strategy for the percentage of swapped patches and denoising steps for each video frame, based on their distance to the condition images. This strategy also introduces position information of condition images into the denoising process, thus making the diffusion model capable of handling multiple condition images at any desired positions.
@@ 4.1 Condition Image Inversion and Replacement
Image Inversion
The overview of our method is illustrated in Fig. 3. First of all, we encode condition images $\bm{x}$ into latent embedding $\tilde{\bm{x}}$ using the pretrained encoder $\mathcal{E}$, e.g., $\tilde{\bm{x}}=\mathcal{E}(\bm{x})=\{\tilde{x}^{(n)}\}_{n=0}^{N-1}$. To fuse the visual condition with T2V denoising process, we invert the clean latent image embedding into noisy representations at each denoising step. We denote the representation at step $t=0$ as $\tilde{\bm{x}}_{0}=\tilde{\bm{x}}$ for consistent notations in diffusion models. The widely-used DDIM inversion [35, 47] is not applicable to our problem because the video diffusion model can not estimate the noise accurately in a single image without temporal information. Instead, we use DDPM-based inversion as an alternative following [34, 37], which directly adds Gaussian noise to $\tilde{\bm{x}}_{0}$ like diffusion process, $\tilde{\bm{x}}_{t}\sim\mathcal{N}\left(\sqrt{\bar{\alpha}_{t}}\;\tilde{\bm{x}}_{0},\sqrt{1-\bar{\alpha}_{t}}\;\bm{I}\right)$. In denoising process, T2V diffusion models typically initialize the noise by drawing a random sample from Gaussian distribution [13]. Recent work has validated reusing noise inverted from images preserves the global structure in image editing [48, 35, 30], and leads to consistent video frames in video generation [37]. Hence, we duplicate the inverted noise $\tilde{\bm{x}}_{T}$ at the final step, and stack them along the temporal dimension as the initial $\bm{z}_{T}$ for video denoising.
Frame Replacement
As shown in Fig. 3, the representation of the $M$ video frames at step $t$ is denoted as $\bm{z}_{t}=\{z_{t}^{(m)}\}_{m=0}^{M-1}$. We leverage the pretrained temporal modeling layer in the T2V diffusion model to propagate visual features of condition images to all video frames. Inspired by [37], we replace the video frame embeddings at the conditioning position $\bm{p}=\{p^{(n)}\}_{n=0}^{N-1}$ with inverted image embeddings, which is formulated as
$FrameReplace+\left(\tilde{\bm{x}}_{t},\bm{z}_{t},\bm{p}\right)=\left\{z_{t}^{(0)},z_{t}^{(1)},\cdots,z_{t}^{(p^{(n)}-1)},\tilde{x}_{t}^{(p^{(n)})},z_{t}^{(p^{(n)}+1)},\cdots,z_{t}^{(M-2)},z_{t}^{(M-1)}\right\},$ (3)
@@ 4.2 Random Patch Swapping
Random patch swapping is essentially replacing the element of $m$-th video frame at position $(i,j)$ with the element of $n$-th condition image at the same position. Each element in the latent space corresponds to a patch in RGB space depending on the spatial downsampling ratio. Then RPS can be formulated as updating the values in $\hat{z}_{t}^{(m)}$ following this equation:
$\hat{z}_{t}^{(m)}[i,j]=\mathcal{M}_{t}^{(m,n)}[i,j]\cdot\tilde{x}_{t}^{(n)}[i,j]+\left(1-\mathcal{M}_{t}^{(m,n)}[i,j]\right)\cdot\hat{z}_{t}^{(m)}[i,j].$ (4)
In this way, the visual cues from condition images are delivered to video frames through swapped patches, and then harmonized with video features by the pretrained diffusion foundation models. Each video frame only swaps patches with the two bounded condition images (the closest condition images before and after it). If all condition images are on the same side of a video frame, this frame only swaps patches with one condition image. We find a large percentage of swapped patches may include overly strong spatial alignment prior, while a small percentage contains insufficient visual features for effective guidance. To make a trade-off, we dynamically reduce the percentage and steps of RPS when the video frame is far from the position of condition image. Thus the percentage $P(m,n,t)$ is written as
$P(m,n,t)=\left\{\begin{array}[]{ll}P_{0}-\delta_{1}\cdot|m-p^{(n)}|,&0\leq t\leq\tilde{t}\\ 0,&\tilde{t}<t\leq T,\end{array}\right.\quad\mathrm{where}\quad\tilde{t}=t_{0}-\delta_{2}\cdot|m-p^{(n)}|.$ (5)
$P_{0}$ and $t_{0}$ are initial swapping percentage and steps, and $\delta_{1}$, $\delta_{2}$ are decreasing strides. Our dynamic control strategy is based on the intuitive hypothesis that video frames around condition images are also perceptually similar, thus demanding more visual cues by patch swapping. Likewise, for frames far from condition images, we reduce constraints of visual conditioning accordingly to encourage more creativity in motion generation. This method also introduces position information of images into a T2V model, which makes our approach applicable to any number of images at any positions.
@@ 5.1 Dataset and Metrics
We run experiments on the widely-used UCF-101 dataset [46] following [14, 8, 37], which is composed of 13,320 videos of 101 action categories. We select 42 action classes including complex actions with high dynamics (e.g., surfing) and small motion with still background (e.g., typing). For each class, we randomly select 25 instances (totaling 1,050 videos) as the test set. We manually rewrite the textual prompts for each action. More details of data preprocessing and textual prompts are described in the supplementary.
Metrics
@@
We use three CLIP-based metrics including (1) CLIP frame-text similarity (CLIP-T) measuring the fidelity of generated frames to textual prompts, (2) CLIP frame similarity (CLIP-F) measuring the consistency across all video frames, and (3) CLIP image similarity (CLIP-I) measuring the agreement between condition images and video frames at other positions. We also use (4) Frechet Video Distance (FVD) [49] to assess the overall video quality, its variant (5) text-conditioned FVD (tFVD) [37] to assess the video-text alignment, and (6) user study for human subjective evaluation.
Implementation Details
@@
For a fair comparison with [37], we implement our FlexTI2V method on ModelScopeT2V [52], which shows great robutness to out-domain T2V generation. We set the number of video frames as $M=16$, and total denoising steps $T=20$ with deterministic DDIM sampling (i.e., $\sigma_{t}=0$ in Eq. 2). The initial percentage and steps of swapping are set as $P_{0}=0.3$ and $t_{0}=10$. The decreasing strides are $\delta_{1}=5\times 10^{-3}$ and $\delta_{2}=0.3$. All experiments are conducted on 1 NVIDIA H100 GPU. Please refer to the supplementary for more implementation details.
Methods Image Animation Methods Video Frame Interpolation
[…]
DC [54] 210.77 832.38 29.86 89.10 80.95 26.20 DC [54] 310.74 731.90 29.69 84.26 78.68 16.20
TI2V-Zero [37] 189.49 586.54 29.84 88.01 82.29 28.40 TRF [7] 247.88 631.54 29.47 91.99 83.35 32.20
FlexTI2V 125.49 558.29 30.97 94.40 84.38 45.40 FlexTI2V 98.41 406.30 30.92 92.34 86.88 51.60
Methods Outpainting Methods Rewinding
[…]
TI2V-Zero [37] 216.80 641.07 28.80 89.01 85.01 38.60 TI2V-Zero [37] 199.01 611.60 29.77 89.21 81.02 39.80
FlexTI2V 167.46 632.71 30.75 95.89 88.34 61.40 FlexTI2V 134.98 596.96 30.95 94.58 85.11 60.20
Table 1: Quantitative comparison with previous methods. “DC” is short for DynamiCrafter. “User” denotes the user study results. $\downarrow$ means a lower score in this metric indicates a better performance. The best results are highlighted with boldface. The orange row refers to our FlexTI2V method.
5.2 Quantitative Comparison with Prior Approaches
We compare with prior approaches quantitatively under three classic settings: (1) image animation – one condition image as the first video frame ($N=0$, $\bm{p}=\{0\}$), (2) rewinding – one condition image as the last video frame ($N=0$, $\bm{p}=\{15\}$), and (3) video frame interpolation – two condition images as the first and last frames ($N=2$, $\bm{p}=\{0,15\}$). We also introduce a new setting, video frame outpainting – one condition image as the middle video frame ($N=1$, $\bm{p}=\{7\}$). We mainly compare with two state-of-the-art training-free methods – TI2V-Zero [37] and time reversal fusion (TRF) [7]. TI2V-Zero is initially designed only for image animation task. We adapt it to outpainting and rewinding tasks by adjusting the order of frame synthesis. Moreover, we also include DynamiCrafter [54] as a training-based baseline for a thorough comparison. More details of baseline implementation are elaborated in the supplementary. The results are shown in Tab. 1.
Single Image Conditioning Tasks
In image animation task, our method outperforms DynamiCrafter and TI2V-Zero by 85.28 and 64.00 respectively in FVD, which suggests an overall higher quality of videos synthesized by our method. In fine-grained evaluation metrics, DynamiCrafter is comparable with TI2V-Zero, but lags behind in tFVD. The possible explanation is that tFVD measures FVD score between generated videos and real videos with the same textual prompts. This is an indirect assessment of video-text semantic alignment which may be more sensitive to distortion and blur. Our FlexTI2V surpasses the best baseline by 28.25 in tFVD and 1.11% in CLIP-T, indicating the better prompt-following capability. In addition, the two baselines have low scores in CLIP-F and CLIP-I, indicating the inconsistency across generated video frames and low fidelity to the condition images. The reason could be that DynamiCrafter is a training-based model that still suffers from the bottleneck of generalization to out-of-domain data. Though TI2V-Zero is a training-free method that uses the same diffusion foundation model as FlexTI2V, it synthesizes the future video frames autoregressively, resulting in compounding errors and cumulative deviations from visual conditions, which can be observed in visualization in Sec. 5.3. In contrast, our method generates all video frames simultaneously in a single denoising process, thus achieving the best performance in CLIP-F and CLIP-I. The prominent improvement (5.30% and 3.43%) also validates the effectiveness of our random patch swapping strategy in incorporating visual features into T2V models. In addition to image animation, we also adapt TI2V-Zero to outpainting and rewinding settings for quantitative comparison. The notable gap further suggests the high flexibility and strong generalization capability of FlexTI2V to image conditioning at any positions.
Multi-Image Conditioning Tasks
In terms of video frame interpolation, DynamiCrafter still lags behind TRF and FlexTI2V due to domain shift. In contrast, TRF shows better performance for this task (measured by FVD) relying on the strong generalization capability of stable video diffusion [2], which further validates the advantage of training-tree method in controllable video generation. Though TRF achieves a high score in CLIP-F and CLIP-I, it still has a limited performance in video-text alignment (tFVD and CLIP-T) due to lack of semantic control by textual prompts. In contrast, our method outperforms both competitors in all the metrics. The significant improvement further validates the flexibility of our method to multi-image conditioning.
User Study
Besides automatic metrics. We also conduct user study for a thorough evaluation. We sample 100 instances from our test set for each task, with each instance evaluated by five raters. Each rater is asked to select the video with best alignment with the textual prompt and condition images. The human preference percentage of each method is presented in Tab. 1. FlexTI2V is selected as the best in the highest percentage of samples compared with other competitors. The results suggest that the videos synthesized by our method aligns better with human’s subjective perception.
@@ 5.2 Quantitative Comparison with Prior Approaches
Figure 4: Comparison with prior methods. The images with blue and pink edges are condition images for each setting, and images with green edges are generated video frames. Our approach synthesizes videos with higher frame consistency and fidelity to condition images than other baseline models.
5.3 Visualization for Qualitative Comparison
Besides quantitative evaluation in Sec. 5.2, we also visualize the videos synthesized by our method and other baselines for a qualitative comparison. As illustrated in Fig. 4, we observe some distortion in the videos generated by DynamiCrafter (e.g., the person in image animation and the knife in video frame interpolation), which may be the explanation for the low performance in Tab. 1. TRF is good at preserving visual features of the two bounded images. However, the video frames synthesized by TRF are blurry around the knife. We speculate that TRF is initially proposed for smooth motion generation and camera viewpoint transition. It may be weak in modeling object state change, such as the onion being cut into small pieces in this example. TI2V-Zero can generate videos with large motion because of the autoregressive generation mechanism. However, the cumulative mistakes weaken the controllability of visual conditions, leading to a low consistency in the frames far from the condition image. In contrast, our method can follow the textual prompts and visual conditions faithfully, and synthesize smooth motions in all TI2V settings. Please refer to the supplementary for more visualization of our method.

### v2

@@ 4 FlexTI2V
The setting of the proposed generic text-image-to-video (TI2V) generation problem is illustrated in Fig. 2. The input includes a textual prompt $y$, NN images $\bm{x}=\{x^{(n)}\}_{n=0}^{N-1}$ as visual conditions, and the expected position index $\bm{p}=\{p^{(n)}\}_{n=0}^{N-1}$ of each image in the output video. The goal is to synthesize a video consisting of $M$ frames $\bm{v}=\{v^{(m)}\}_{m=0}^{M-1}$ with each condition image at the expected position (i.e., $v^{(p^{(n)})}=x^{(n)}$ where $p^{(n)}\in[0,M-1]$). The output video is supposed to show a reasonable transition across condition images following the textual prompt. Our problem setting unifies many existing TI2V tasks in one formulation, such as image animation ($N=1$, $\bm{p}=\{0\}$), video rewinding ($N=1$, $\bm{p}=\{M-1\}$) and video frame interpolation ($N=2$, $\bm{p}=\{0,M-1\}$).
The major challenge of training-free TI2V generation is harmonizing the distinct denoising directions controlled by texts and images solely relying on frozen model parameters. The experiments of MAE [he2022masked] validate that global image information can be carried by and inferred from a portion of local patches. Inspired by this finding, the key insight of our FlexTI2V method is contextualizing each synthetic video frame with visual conditioning, by randomly swapping a portion of patches between video frames and condition images in the latent space at early denoising steps. To balance the fidelity to condition images and creativity in motion generation, we use a dynamic control strategy for the percentage of swapped patches and denoising steps for each video frame, based on their distance to the condition images. This strategy also introduces position information of condition images into the denoising process, thus making the diffusion model capable of handling multiple condition images at any desired positions.
@@ 4.1 Condition Image Inversion and Replacement
Image Inversion The overview of our method is illustrated in Fig. 3. First of all, we encode condition images $\bm{x}$ into latent embedding $\tilde{\bm{x}}$ using the pretrained encoder $\mathcal{E}$, e.g., $\tilde{\bm{x}}=\mathcal{E}(\bm{x})=\{\tilde{x}^{(n)}\}_{n=0}^{N-1}$. To fuse the visual condition with T2V denoising process, we invert the clean latent image embedding into noisy representations at each denoising step. We denote the representation at step $t=0$ as $\tilde{\bm{x}}_{0}=\tilde{\bm{x}}$ for consistent notations in diffusion models. The widely-used DDIM inversion [mokady2023null, titov2024guide] is not applicable to our problem because the video diffusion model can not estimate the noise accurately in a single image without temporal information. Instead, we use DDPM-based inversion as an alternative following [meng2022sdedit, ni2024ti2v], which directly adds Gaussian noise to $\tilde{\bm{x}}_{0}$ like diffusion process, $\tilde{\bm{x}}_{t}\sim\mathcal{N}\left(\sqrt{\bar{\alpha}_{t}}\;\tilde{\bm{x}}_{0},\ (1-\bar{\alpha}_{t})\bm{I}\right)$. In denoising process, T2V diffusion models typically initialize the noise by drawing a random sample from Gaussian distribution [ho2020denoising]. Recent work has validated reusing noise inverted from images preserves the global structure in image editing [tumanyan2023plug, mokady2023null, liu2024towards], and leads to consistent video frames in video generation [ni2024ti2v]. Hence, we duplicate the inverted noise $\tilde{\bm{x}}_{T}$ at the final step, and stack them along the temporal dimension as the initial $\bm{z}_{T}$ for video denoising.
Frame Replacement As shown in Fig. 3, the representation of the $M$ video frames at step $t$ is denoted as $\bm{z}_{t}=\{z_{t}^{(m)}\}_{m=0}^{M-1}$. We leverage the pretrained temporal modeling layer in the T2V diffusion model to propagate visual features of condition images to all video frames. Inspired by [ni2024ti2v], we replace the video frame embeddings at the conditioning position $\bm{p}=\{p^{(n)}\}_{n=0}^{N-1}$ with inverted image embeddings, which is formulated as
$\verb+FrameReplace+\left(\tilde{\bm{x}}_{t},\bm{z}_{t},\bm{p}\right)=$ (3)
$\left\{z_{t}^{(0)},z_{t}^{(1)},\cdots,z_{t}^{(p^{(n)}-1)},\tilde{x}_{t}^{(p^{(n)})},z_{t}^{(p^{(n)}+1)},\cdots,z_{t}^{(M-2)},z_{t}^{(M-1)}\right\},$
@@ 4.2 Random Patch Swapping
Methods Image Animation Methods Video Frame Interpolation
[…]
DC [xing2024dynamicrafter] 26.20 210.77 832.38 29.86 89.10 80.95 DC [xing2024dynamicrafter] 16.20 310.74 731.90 29.69 84.26 78.68
TI2V-Zero [ni2024ti2v] 28.40 189.49 586.54 29.84 88.01 82.29 TRF [feng2024explorative] 32.20 247.88 631.54 29.47 91.99 83.35
FlexTI2V 45.40 125.49 558.29 30.97 94.40 84.38 FlexTI2V 51.60 98.41 406.30 30.92 92.34 86.88
Methods Outpainting Methods Rewinding
[…]
TI2V-Zero [ni2024ti2v] 38.60 216.80 641.07 28.80 89.01 85.01 TI2V-Zero [ni2024ti2v] 39.80 199.01 611.60 29.77 89.21 81.02
FlexTI2V 61.40 167.46 632.71 30.75 95.89 88.34 FlexTI2V 60.20 134.98 596.96 30.95 94.58 85.11
Table 1: Quantitative comparison with previous methods. “DC” is short for DynamiCrafter. “User” denotes the user study results. $\downarrow$ means a lower score in this metric indicates a better performance. The orange row refers to our method.
DC [xing2024dynamicrafter] TI2V-Zero [ni2024ti2v] TRF [feng2024explorative] FlexTI2V
Inference Time (s) 9.45 275.14 60.05 2.71
Table 2: Comparison of model efficiency in inference. “DC” is short for DynamiCrafter. The orange column refers to our method.
Random patch swapping is essentially replacing the element of $m$-th video frame at position $(i,j)$ with the element of $n$-th condition image at the same position. Each element in the latent space corresponds to a patch in RGB space. Then RPS can be formulated as updating the values in $\hat{z}_{t}^{(m)}$ following this equation:
$\footnotesize\hat{z}_{t}^{(m)}[i,j]=\mathcal{M}_{t}^{(m,n)}[i,j]\cdot\tilde{x}_{t}^{(n)}[i,j]+\left(1-\mathcal{M}_{t}^{(m,n)}[i,j]\right)\cdot\hat{z}_{t}^{(m)}[i,j].$ (4)
In this way, the visual cues from condition images are delivered to video frames through swapped patches, and then harmonized with video features by the pretrained diffusion foundation models. Each video frame only swaps patches with the two bounded condition images (the closest condition images before and after it). If all condition images are on the same side of a video frame, this frame only swaps patches with one condition image. We find a large percentage of swapped patches may include overly strong spatial alignment prior, while a small percentage contains insufficient visual features for effective guidance. To make a trade-off, we dynamically reduce the percentage and steps of RPS when the video frame is far from the position of condition image. The percentage $P(m,n,t)$ is written as
$P(m,n,t)=$ $\left\{\begin{array}[]{ll}P_{0}-\delta_{1}\cdot|m-p^{(n)}|,&0\leq t\leq\tilde{t}\\ 0,&\tilde{t}<t\leq T,\end{array}\right$. (5)
$\hskip 18.49988pt\mathrm{where}\qquad\tilde{t}=t_{0}-\delta_{2}\cdot|m-p^{(n)}|.$
$P_{0}$ and $t_{0}$ are initial swapping percentage and steps, and $\delta_{1}$, $\delta_{2}$ are decreasing strides. Our dynamic control strategy is based on the intuitive hypothesis that video frames around condition images are also perceptually similar, thus demanding more visual cues by patch swapping. Likewise, for frames far from condition images, we reduce constraints of visual conditioning accordingly to encourage more creativity in motion generation. This method also introduces position information of images into a T2V model, which makes our approach applicable to any number of images at any positions. The entire algorithm formulation is shown in Algorithm A in the supplementary.
@@ 5.1 Dataset and Metrics
Dataset We run experiments on the widely-used UCF-101 dataset [soomro2012ucf101] following [ho2022video, fu2023tell, ni2024ti2v], which is composed of 13,320 videos of 101 action categories. We select 42 action classes including complex actions with high dynamics (e.g., surfing) and small motion with still background (e.g., typing). For each class, we randomly select 25 instances (totaling 1,050 videos) as the test set. We manually rewrite the textual prompts for each action. Please also refer to Supp. Sec. B.2 for experiments on another synthetic dataset.
@@
Metrics We use three CLIP-based metrics including (1) CLIP frame-text similarity (CLIP-T) measuring the fidelity of generated frames to textual prompts, (2) CLIP frame similarity (CLIP-F) measuring the consistency across all video frames, and (3) CLIP image similarity (CLIP-I) measuring the agreement between condition images and video frames at other positions. We also use (4) Frechet Video Distance (FVD) [unterthiner2018towards] to assess the overall video quality, its variant (5) text-conditioned FVD (tFVD) [ni2024ti2v] to assess the video-text alignment, and (6) user study for human evaluation.
@@
Implementation Details For a fair comparison with [ni2024ti2v], we implement our FlexTI2V method on ModelScopeT2V [wang2023modelscope], which shows great robutness to out-domain T2V generation. We set the number of video frames as $M=16$, and total denoising steps $T=20$ with deterministic DDIM sampling (i.e., $\sigma_{t}=0$ in Eq. 2). The initial percentage and steps of swapping are set as $P_{0}=0.3$ and $t_{0}=10$. The decreasing strides are $\delta_{1}=5\times 10^{-3}$ and $\delta_{2}=0.3$. All experiments are conducted on 1 NVIDIA H100 GPU. Please refer to Supp. Sec. C.2 for more implementation details.
@@
Figure 4: Comparison with prior methods. The images with blue and pink edges are condition images for each setting, and images with green edges are generated video frames. Our approach synthesizes videos with higher frame consistency and fidelity to condition images than other baseline models in various settings.
Figure 5: Generated videos of our method implemented to Wan2.1-T2V. The images with blue edges are condition images and images with green edges are generated video frames. Please refer to the supplementary for corresponding videos and extra demos.
5.2 Quantitative Comparison
We compare with prior approaches quantitatively under three classic settings: (1) image animation – one condition image as the first video frame ($N=0$, $\bm{p}=\{0\}$), (2) rewinding – one condition image as the last video frame ($N=0$, $\bm{p}=\{15\}$), and (3) video frame interpolation – two condition images as the first and last frames ($N=2$, $\bm{p}=\{0,15\}$). We also introduce a new setting, video frame outpainting – one condition image as the middle video frame ($N=1$, $\bm{p}=\{7\}$). We mainly compare with two state-of-the-art training-free methods – TI2V-Zero [ni2024ti2v] and time reversal fusion (TRF) [feng2024explorative]. TI2V-Zero is initially designed only for image animation task. We adapt it to outpainting and rewinding tasks by adjusting the order of frame synthesis. Moreover, we also include DynamiCrafter [xing2024dynamicrafter] as a training-based baseline for a thorough comparison. More details of baseline implementation are elaborated in Supp. Sec. C.3. The results are shown in Tab. 1.
Single-Image Conditioning For image animation, outpainting and rewinding tasks, our method outperforms DynamiCrafter and TI2V-Zero by a notable margin in FVD, which suggests an overall higher quality of videos synthesized by our method. In fine-grained evaluation metrics, our FlexTI2V method surpasses the best baseline by 28.25/8.36/14.64 in tFVD and 1.11%/1.95%/1.18% in CLIP-T for the three tasks, indicating the best prompt-following capability. In addition, the two baselines have low scores in CLIP-F and CLIP-I, indicating the inconsistency across generated video frames and low fidelity to the condition images. The reason could be that DynamiCrafter is a training-based model which still suffers from the bottleneck of generalization to out-of-domain data. Though TI2V-Zero is a training-free method, it synthesizes the future video frames autoregressively, resulting in compounding errors and cumulative deviations from visual conditions, which can also be observed in visualization in Sec. 5.3. In contrast, our method generates all video frames simultaneously in a single denoising process, thus achieving the best performance in CLIP-F and CLIP-I. The prominent improvement (5.30%/6.88%/5.37% and 2.09%/3.33%/4.09%) also validates the effectiveness of our random patch swapping strategy for various image conditioning settings.